# thanks: These authors contribute equallythanks: These authors contribute equally

### v1

@@ Acknowledgements.
The NARMA task setup follows the pioneering work on quantum reservoir computing Fujii and Nakajima (2017), which has also been adopted in several subsequent studies Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024). Two types of NARMA dynamics are considered:
@@
.5 V.    Numerical simulation of 9-spin reservoir computing for NARMA tasks
@@
To validate the reliability of our experimental results, we performed numerical simulations of the nine-spin quantum reservoir computing model on the NARMA task, following the procedure described in Section .3 and using the parameters listed in Table 2. The simulations were implemented using the Python package QuTiP Lambert et al. (2024). For the polarization setting, because the realistic polarization values in room‑temperature NMR (on the order of $10^{-5}$) can lead to numerical instability when solving the Lindblad equation, we scale the carbon and proton polarizations to $1,\,3.98]\times 10^{-4}$ (proportional to their gyromagnetic ratios) to balance numerical robustness with experimental realism.
The simulation results, presented in FIG. 7(a), qualitatively resemble those in Fig. 3 of the main text. Quantitative differences are attributed to deviations between the simulated dynamics and the actual experimental processes, including relaxation mechanisms that extend beyond independent single-spin relaxation and unidentified systematic errors. In our simulations, we incorporated off-resonance effects during the input stage (i.e., imperfect $R_{x}$ pulses) and introduced Gaussian noise during the readout stage, based on an experimental signal-to-noise ratio (SNR) of approximately 300 in the thermally equilibrated spectrum. Spatial field inhomogeneity was also taken into account; however, it did not significantly affect the results and appears to be incorporated as a fixed mechanism of the reservoir. Accurate modeling of open quantum many-body dynamics remains a significant challenge and warrants further investigation.
@@
Figure 7: Prediction performance (evaluated by $R^{2}$) on NARMA tasks with $n$ varying from 2 to 20, obtained from numerical simulations of (a) quantum reservoir computing and (b) classical spin reservoir computing. Red and blue lines represent input intervals of $\tau=$0.3\text{\,}\mathrm{s}$\$ and $0.01\text{\,}\mathrm{s}$, respectively, while solid and dashed lines indicate results obtained with time multiplexing and single-time measurements.
@@
To elucidate the effect of quantumness, we compare our QRC model with the corresponding classical spin reservoir computing model, obtained by treating the spins as classical unit vectors Kornjača et al. (2024). The dynamics of the $i$-th spin are described by
@@
where $\Omega_{i}=2\pi\nu_{i}+2\pi\sum_{j\neq i}J_{ij}S_{z}^{(j)}$ represents the effective instantaneous field exerted on spin-$i$ by the surrounding spins. Hence, the dynamics of $N=9$ classical spins are characterized by $3N=27$ differential equations. The simulation results for the classical spin reservoir, with experimental settings identical to those used for QRC, are shown in FIG. 7(b). These results demonstrate that the optimal performance achievable by the classical spin reservoir remains inferior to that of the quantum reservoir, even when time multiplexing is employed. This is because classical spins lack the additional degrees of freedom, such as quantum correlations, that can be exploited via time multiplexing, thereby highlighting the advantage of quantumness. Moreover, we attribute the improvement achieved by time multiplexing in classical systems to the mitigation of random noise through averaging over multiple readouts and the extraction of information from the nonlinear dynamics induced by $\Omega_{i}S_{x,y}^{(i)}\sim\sum_{j\neq i}S_{z}^{(j)}S_{x,y}^{(i)}$.
.6 VI.    Weather forecasting and echo state networks
@@
The weather dataset referenced in the main text is sourced from Kaggle dai (2019), a prominent online platform for data science competitions and resources. The dataset provides daily weather data for Delhi, India, covering 1,575 days from January 1, 2013, to April 24, 2017. It includes four features: temperature, humidity, wind speed, and air pressure. For simplicity of experiments, this study focuses on the first two features, temperature and humidity, whose values are presented in FIG. 8. Direct multi-step forecast is performed for the two-dimensional temporal sequence by training separate reservoir weights to model the mapping from $\{{\bf y}_{k}\}$ to $\{{\bf y}_{k+h}\}$ for different forecast horizons $h$, effectively enabling multitasking Taieb et al. (2012). The dataset is partitioned into 374 steps for washout, 600 steps for training, and the remaining steps for testing. For the testing phase, 600 steps are used for one-step-ahead prediction, as shown in Fig. 4(a) of the main text, while 556 steps are used for multi-step-ahead prediction with a maximum forecast horizon of $h_{\max}=45$.
@@
The baseline model employs Echo State Networks (ESNs), a widely used reservoir computing model based on recurrent neural networks Jaeger and Haas (2004). For an ESN with $M$ network nodes (denoted as ${\rm ESN}(M)$), the state update and output are described by
@@
We implemented the ESNs using the Python package reservoirpy Trouvain et al. (2020), where ${\bf W}_{\rm in}$ is sampled from a Bernoulli distribution over $\{-1,1\}$ and $\bf W$} is sampled from a standard normal distribution. The connectivity $k$ and spectral radius $r$ of of the recurrent weight matrix $\bf W$}, defined as the fraction of nonzero entries and the largest absolute eigenvalue, respectively, are closely related to the computational power of ESNs. In this work, these hyperparameters were determined via grid search, yielding $(k,r)$ pairs of (0.025, 0.99), (0.01, 0.99), (0.0025, 0.99), and (0.0025, 0.99) for ESN(500), ESN(1000), ESN(5000), and ESN(10000), respectively. FIG. 11 presents the 45-day-ahead prediction results of different models.
We evaluated support vector regression (SVR) with a radial basis function (RBF) kernel as a nonlinear post–processor for the reservoir readouts. Our implementation employed the SVR module from the scikit-learn library Pedregosa et al. (2011), with the regularization parameter $C$ and the tolerance margin $\epsilon$ systematically tuned. Despite exploring a wide range of hyperparameter settings, SVR did not deliver significant performance improvements for ESNs. FIG. 9 reports the best results obtained.
@@
.7 VII.    Effects of $T_{1}$ and $T_{2}$ relaxation on the fading memory in quantum reservoir computing
@@
In this section, we examine the influence of the $T_{1}$ and $T_{2}$ relaxation on the fading memory of the quantum reservoir, as quantified by the short-term memory (STM) capacity, through numerical simulations. STM capacity quantifies the ability of the reservoir to reconstruct past inputs from its current state, which is a fundamental property for processing temporal information.
@@
First, we reiterate the necessity of incorporating relaxation processes. As mentioned in the main text, an effective quantum reservoir should possess a high-dimensional state space, nonlinear input–output transformations, and fading memory. In unitary systems (commonly employed in quantum algorithm frameworks) the first two properties emerge naturally: the exponential scaling of the Hilbert space provides a high-dimensional representation, while nonlinearity arises from two mechanisms: (i) the input encoding (e.g., rotations yielding sine and cosine responses) Govia et al. (2022) and (ii) the multiplicative interaction between sequential inputs Fujii and Nakajima (2017); Pérez-Salinas et al. (2020). In contrast, fading memory requires that, after a sufficiently long input sequence, the reservoir state converges to a unique trajectory independent of its initial conditions. Due to the distance-preserving nature of unitary dynamics, nonunitary processes (such as $T_{1}$ and $T_{2}$ relaxation) are essential for gradually erasing old information.
@@
Table 3: Parameters of the three-spin diethyl fluoromalonate sample. The left portion presents chemical shifts (diagonal elements) and $J$-couplings (off-diagonal elements), while the right lists the corresponding relaxation times.
In our numerical simulations, we employ a three‑spin system with parameters drawn from diethyl fluoromalonate, a widely used NMR molecular sample, as detailed in Table 3. The spin polarizations are set to $1,\,3.98,\,3.74]\times 10^{-4}$ for carbon, proton, and fluorine respectively. Considering that the input interval affects the memory capacity, evolution duration $\tau$ is varied from $10^{-4}~{\rm s}$ to $10^{1}~{\rm s}$ to explore its influence. To avoid the uncertainty caused by the selection of partial measurements, the readout is obtained by measuring $4^{3}-1=63$ Pauli operators to capture the full state information. These expectation values are then recalibrated by dividing by the polarization and further perturbed by adding Gaussian noise with $\sigma=10^{-5}$ to improve the robustness of linear regression. For the dataset, we generate a prolonged random sequence $\{s_{k}\}$ uniformly distributed in $[0,1]$. The training and test sets are taken from the tail of the sequence, with sizes of 3000 and 1000 points, respectively. A washout period of size $\max\left\{\left\lceil 10/\tau\right\rceil,t_{d}^{\max}\right\}$ is applied immediately prior to the training data to eliminate the effects of the initial state, lasting approximately 10 seconds.
@@
FIG. 10(a) illustrates the total STM capacity $C_{\rm STM}$ for various configurations of input intervals and relaxation mechanisms. As expected, $C_{\rm STM}$ vanishes under unitary evolution due to the absence of fading memory. In a unitary system, the reservoir state retains the full influence of the initial state and all prior inputs, preventing the selective extraction of relevant recent information. However, although the introduction of $T_{2}$ relaxation (phase damping) leads to some information erasure, it alone is insufficient to achieve a meaningful fading memory. Under unitary dynamics, the quantum system continuously interconverts population and coherence, while dephasing gradually eliminates coherence. This interplay ultimately depolarizes the reservoir, driving it toward the trivial maximally mixed state $\mathds{1}/2^{N}$.
@@
Thus, not every dissipative process is suitable for quantum reservoir computing; the key requirement is a process that drives the system toward a nontrivial state—a role fulfilled by $T_{1}$ relaxation (generalized amplitude damping). Several studies have underscored the necessity of non-unital channels (i.e., channels that do not preserve the identity) in quantum reservoir computing Mart´ınez-Peña and Ortega (2023); Monzani et al. (2024). From an information-theoretic perspective, we conjecture that entropy-decreasing processes are essential for achieving effective fading memory. In amplitude damping channels, the system is driven toward an equilibrium state with lower entropy than the maximally mixed state. Moreover, previous schemes employing input qubit reset Fujii and Nakajima (2017) and probabilistic state reset Chen et al. (2020) incorporate entropy-reducing state preparation processes. In contrast, unitary operations preserve entropy, whereas dephasing increases it Nielsen and Chuang (2001).
@@
The configuration incorporating $T_{1}$ relaxation exhibits a nonzero STM capacity, with $C_{\rm STM}$ initially increasing and then decreasing as the input interval varies. FIG. 10(b) and (c) show the corresponding behavior of $C(t_{d})$ as a function of delay. In the long-time limit ($\tau\gg T_{1}$), the reservoir relaxes to an equilibrium state independent of the inputs, causing $C_{\rm STM}$ to vanish. For relatively large $\tau$, relaxation erases most of the input history while only retains the most recent inputs, manifested as a narrow plateau ($C(t_{d})\approx 1$) followed by a rapid decline in $C(t_{d})$. As $\tau$ decreases, the reservoir incorporates information from more input steps; however, nonlinear mixing of these inputs can reduce the effectiveness of linear memory retrieval, resulting in an initial $C(0)<1$ and a gradual decay to zero at larger $t_{d}$. With further decrease in $\tau$, the overall values of $C(t_{d})$ diminish, and in the short-$\tau$ limit, the system behaves similarly to a unitary system, lacking effective fading memory. These results suggest the existence of an optimal input interval that balances the number of retained inputs with high memory fidelity at each delay.

### v2

@@ Appendix D The NARMA task setup and experimental prediction results
The NARMA task setup follows the pioneering work on quantum reservoir computing [19], which has also been adopted in several subsequent studies [77, 86, 55]. Two types of NARMA dynamics are considered:
@@
Appendix E Numerical characterization of spin-based quantum reservoir dynamics
In this section, we present the numerical implementation of the input injection and the subsequent system evolution in the quantum reservoir. The normalized input $\bar{s}_{k}\in[0,1]$ is encoded as a rotation angle $\theta_{k}=\arcsin(\bar{s}_{k})$ of an RF pulse applied to a specific nuclear channel (e.g., proton), effecting a global $x$-rotation on the corresponding spin set $S_{\rm in}$:
$\prod_{n\in S_{\rm in}}R^{n}_{x}(\theta_{k})=\exp\left[-i\,t_{k}\sum_{n\in S_{\rm in}}\frac{\omega_{d}}{2}\sigma_{x}^{n}\right],$ (11)
with a fixed driving frequency $\omega_{d}$ and an adjusted pulse duration $t_{k}$ satisfying $\omega_{d}t_{k}=\theta_{k}$. More precisely, when the internal Hamiltonian evolution during the pulse duration is taken into account, the actual unitary operation becomes
$U_{s_{k}}=\exp\!\left[-i\,t_{k}\left(\sum_{n\in S_{\rm in}}\frac{\omega_{d}}{2}\sigma_{x}^{n}+H\right)\right],$ (12)
where the system Hamiltonian is
This introduces a slight deviation of the effective rotation axis from the $x$ direction.
The system evolution is governed by the Lindblad master equation:
[…]
where $\mathcal{R}$ denotes the relaxation channel. For a single spin, the relaxation channel is given by
with the collapse operators defined as
$\rho_{\rm eq}=\frac{1}{2}\begin{bmatrix}1+p&0\\ 0&1-p\end{bmatrix}$, (17)
at a rate $T_{1}^{-1}$, where $p$ represents the spin polarization (with $p=1$ recovering the standard amplitude damping channel) [61]. The operator $L_{z}$ accounts for pure dephasing due to spin-spin interactions and field inhomogeneities, with a characteristic time $T_{\phi}$. Combined with amplitude damping, the effective decoherence rate is given by
$T_{2}^{-1}=\left(2T_{1}\right)^{-1}+T_{\phi}^{-1}.$ (18)
$\mathcal{R}[\rho]=\sum_{n=1}^{N}\mathcal{R}^{(n)}[\rho],$ (19)
@@ Appendix E Numerical characterization of spin-based quantum reservoir dynamics
Appendix F Numerical simulation of 9-spin reservoir computing for NARMA tasks
@@
To validate the reliability of our experimental results, we performed numerical simulations of the nine-spin quantum reservoir computing model on the NARMA task, following the procedure described in Section E and using the parameters listed in Table 2. The simulations were implemented using the Python package QuTiP [42]. For the polarization setting, because the realistic polarization values in room‑temperature NMR (on the order of $10^{-5}$) can lead to numerical instability when solving the Lindblad equation, we scale the carbon and proton polarizations to $1,\,3.98]\times 10^{-4}$ (proportional to their gyromagnetic ratios) to balance numerical robustness with experimental realism.
The simulation results, presented in Fig. 7(a), are qualitatively consistent with the experimental results in Fig. 3 of the main text. Quantitative differences are attributed to deviations between the simulated dynamics and the actual experimental processes, including cross-correlated relaxation [39] and unidentified systematic errors. A detailed analysis of experimental imperfections and their effects on QRC performance is provided in Section G. Accurate modeling of open quantum many-body dynamics remains a significant challenge and warrants further investigation.
@@
Figure 7: Prediction performance (evaluated by $R^{2}$) on NARMA tasks with $n$ varying from 2 to 20, obtained from numerical simulations of (a) quantum spin reservoir and (b) classical spin reservoir. Red and blue lines represent input intervals of $\tau=$0.3\text{\,}\mathrm{s}$\$ and $0.01\text{\,}\mathrm{s}$, respectively, while solid and dashed lines indicate results obtained with time multiplexing and single-time measurements.
@@
To elucidate the effect of quantumness, we compare our QRC model with the corresponding classical spin reservoir computing model, obtained by treating the spins as classical unit vectors [38]. The dynamics of the $i$-th spin are described by
@@
where $\Omega_{i}=2\pi\nu_{i}+2\pi\sum_{j\neq i}J_{ij}S_{z}^{(j)}$ represents the effective instantaneous field exerted on spin-$i$ by the surrounding spins. Hence, the dynamics of $N=9$ classical spins are characterized by $3N=27$ differential equations. The simulation results for the classical spin reservoir, with experimental settings identical to those used for QRC, are shown in FIG. 7(b). These results demonstrate that the optimal performance achievable by the classical spin reservoir remains inferior to that of the quantum reservoir, even when time multiplexing is employed. This is because classical spins lack the additional degrees of freedom, such as quantum correlations, that can be exploited via time multiplexing, thereby highlighting the advantage of quantumness.
Appendix G Robustness analysis against experimental imperfections
In this section, we analyze the robustness of QRC against different experimental imperfections. The impact of such imperfections mainly depends on whether they are time-invariant or time-varying.
  • (a)
    Time-invariant systematic deviations can be regarded as part of the intrinsic reservoir dynamics, effectively transforming the ideal quantum map into a modified yet still deterministic one, whose influence is not necessarily positive or negative. In particular, constant linear offsets in the readout can be fully absorbed by linear regression during training, yielding the same prediction accuracy as in the ideal case.
  • (b)
    Time-varying or random errors disrupt the deterministic nature of the reservoir mapping, causing identical input sequences to evolve into inconsistent trajectories over time. This randomness blurs the state-space representation, reduces the fidelity of information retrieval, and thus weakens the reservoir’s separability, memory capacity, and predictive accuracy.
To quantify these effects, we numerically examined three representative imperfections: (a) pulse errors, (b) relaxation-time drift, and (c) measurement noise. To reduce computational cost, simulations were performed on a six-spin subsystem ($\mathrm{C}_{1}$$\mathrm{C}_{3}$ and $\mathrm{H}_{1}$$\mathrm{H}_{3}$) of the nine-spin molecule, with all other parameters identical to those used in experiments. The input interval was fixed at $\tau=0.01\,\mathrm{s}$, and time-multiplexed readout was employed. The performance was evaluated using the average prediction accuracy $\overline{R^{2}}$ across NARMA2–NARMA20 tasks.
Figure 8: Average NARMA performance (evaluated by $\overline{R^{2}}$) as a function of error scale of (a) pulse strength, (b) relaxation time drift, and (c) measurement noise. Simulations use a six-spin subsystem ($\mathrm{C}_{1}$$\mathrm{C}_{3}$ and $\mathrm{H}_{1}$$\mathrm{H}_{3}$) of the nine-spin molecule, with input interval $\tau=0.01\,\mathrm{s}$ and time-multiplexed readout. Each point represents the average over 20 random realizations, with the standard deviation shown as the error bar. The dashed line represents the ideal case.
Pulse imperfections. One relevant source is the finite-duration effect of RF pulses: during the pulse, the system evolves under both the control field and the internal Hamiltonian as described in Eq. (12), resulting in a time-invariant deviation from an ideal rotation. In our setting, this induces a slight tilt of the effective rotation axis from the $x$ direction, and changes $\overline{R^{2}}$ by only 0.001, indicating a negligible impact on overall performance.
Another source is the deviation of RF amplitude, modeled as
$U_{s_{k}}^{\epsilon}=\exp\!\left[-i\,t_{k}\left(\sum_{n\in S_{\rm in}}\frac{\omega_{d}}{2}(1+\epsilon)\sigma_{x}^{n}+H\right)\right],$ (21)
where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Two cases are considered: (i) time-invariant offsets arising from spatial RF inhomogeneity and (ii) time-varying fluctuations representing stochastic drift across different pulses within a single experiment. As shown in Fig. 8(a), time-invariant offsets (blue curve) have almost no influence on $\overline{R^{2}}$, indicating that such systematic deviations are effectively incorporated into the reservoir dynamics without disrupting its operation, demonstrating the intrinsic robustness of QRC. (Note that some static deviations can still degrade performance, such as strong decoherence.) In contrast, time-varying errors (red curve) lead to a monotonic decline in $\overline{R^{2}}$ with increasing error scale, decreasing by about 0.004 at $\sigma=10^{-2}$ and nearly 0.15 at $\sigma=10^{-1}$. In our experiments, the dominant pulse imperfections arise from static RF field inhomogeneity across the sample volume, typically up to 5% depending on the probe and sample geometry, whereas temporal pulse-to-pulse amplitude fluctuations are generally negligible.
Relaxation time drift. Relaxation times may drift over long experimental runs or fluctuate due to environmental variations. We model this effect as
$T^{\epsilon}_{1}=T_{1}^{0}(1+\epsilon_{1}),\quad T^{\epsilon}_{2}=T_{2}^{0}(1+\epsilon_{2}),$ (22)
with $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ updated every 100 steps to emulate slow fluctuations. As shown in Fig. 8(b), $\overline{R^{2}}$ decreases by 0.003 at $\sigma=10^{-2}$ and by 0.073 at $\sigma=10^{-1}$, indicating strong robustness against small quasi-static drift, while large fluctuations reduce performance due to loss of dynamical determinism (a drop of 0.513 at $\sigma=10^{0}$). In NMR experiments, relaxation time drift may arise from temperature variations, magnetic-field instability, or slow chemical changes in the sample solution. In our setup, however, the longest runs last only about 10 hours, during which variations in $T_{1}$ and $T_{2}$ are negligible under active temperature stabilization.
Furthermore, if the relaxation times can be accurately tuned, they could serve as an additional resource for improving QRC performance. From a dynamical perspective, maintaining $T_{1}$ and $T_{2}$ within a comparable range ensures a feasible operating regime. As discussed in Section I, the two relaxation processes play competing roles: $T_{1}$ relaxation drives the system toward the ground or a thermal-equilibrium state with finite polarization, whereas $T_{2}$ relaxation, together with noncommuting unitary dynamics such as $R_{x}$, depolarizes the system toward a maximally mixed state with vanishing polarization. When $T_{2}\ll T_{1}$, depolarization dominates and the reservoir operates at extremely low polarization, making it practically infeasible because the output signal becomes easily obscured by noise. Conversely, tunable relaxation allows for richer noncommuting dynamics while avoiding excessive depolarization, potentially enabling more expressive information processing. From a task-adaptation perspective, $T_{1}$ and $T_{2}$ determine the reservoir’s effective memory depth. Thus, tuning them can in principle optimize the match between the reservoir memory and the temporal structure of a given task, thereby improving predictive performance [73].
Measurement noise. Measurement errors are modeled as additive Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ applied to the normalized FID signal. Each simulated FID signal was normalized by the average maximum amplitude of all FID signals over the dataset, after which Gaussian noise $\epsilon$ was added. The parameter $\sigma$ therefore serves as an inverse measure of the signal-to-noise ratio (SNR). As shown in Fig. 8(c), random noise monotonically degrades performance, with $\overline{R^{2}}$ decreasing by 0.003, 0.025, 0.157, and 0.563 for $\sigma=10^{-4}$, $10^{-3}$, $10^{-2}$, and $10^{-1}$, respectively. When SNR exceeds $10^{3}$, the impact of measurement noise becomes negligible, consistent with our experimental conditions.
Comparing the three types of imperfections—time-varying pulse errors, relaxation time drift, and measurement noise—we observe that measurement noise is the dominant source of the degradation of QRC performance, as it leads to comparable drops in $\overline{R^{2}}$ at smaller error scales $\sigma$. This may arise because dynamical imperfections are partially mitigated by the reservoir’s fading memory mechanism, whereas measurement noise directly affects the readout and cannot be compensated by subsequent processing. These findings provide practical guidance for improving experimental implementations of QRC. In addition, we suggest that system stability can be promoted not only by improving experimental hardware but also by leveraging the native operations and intrinsic dynamics of the physical platform, enabling simpler, faster, and more reliable realizations of QRC.
Appendix H Weather forecasting and echo state networks
@@
The weather dataset referenced in the main text is sourced from Kaggle [15], a prominent online platform for data science competitions and resources. The dataset provides daily weather data for Delhi, India, covering 1,575 days from January 1, 2013, to April 24, 2017. It includes four features: temperature, humidity, wind speed, and air pressure. For simplicity of experiments, this study focuses on the first two features, temperature and humidity, whose values are presented in FIG. 9. Direct multi-step forecast is performed for the two-dimensional temporal sequence by training separate reservoir weights to model the mapping from $\{{\bf y}_{k}\}$ to $\{{\bf y}_{k+h}\}$ for different forecast horizons $h$, effectively enabling multitasking [78]. The dataset is partitioned into 374 steps for washout, 600 steps for training, and the remaining steps for testing. For the testing phase, 600 steps are used for one-step-ahead prediction, as shown in Fig. 4(a) of the main text, while 556 steps are used for multi-step-ahead prediction with a maximum forecast horizon of $h_{\max}=45$.
@@
The baseline model employs Echo State Networks (ESNs), a widely used reservoir computing model based on recurrent neural networks [28]. For an ESN with $M$ network nodes (denoted as ${\rm ESN}(M)$), the state update and output are described by
@@
We implemented the ESNs using the Python package reservoirpy [81], where ${\bf W}_{\rm in}$ is sampled from a Bernoulli distribution over $\{-1,1\}$ and $\bf W$} is sampled from a standard normal distribution. The connectivity $k$ and spectral radius $r$ of the recurrent weight matrix $\bf W$}, defined as the fraction of nonzero entries and the largest absolute eigenvalue, respectively, are closely related to the computational power of ESNs. In this work, these hyperparameters were determined via grid search, yielding $(k,r)$ pairs of (0.025, 0.99), (0.01, 0.99), (0.0025, 0.99), and (0.0025, 0.99) for ESN(500), ESN(1000), ESN(5000), and ESN(10000), respectively. FIG. 11 presents the 45-day-ahead prediction results of different models.
We note that for both QRC and ESNs, the prediction accuracy for humidity is generally lower than that for temperature. This difference mainly reflects their distinct statistical and physical characteristics [40, 72]. Temperature evolution is primarily governed by large-scale, slowly varying processes such as diurnal and seasonal cycles and long-term ocean–atmosphere coupling, resulting in smooth, low-frequency dynamics with strong temporal correlations that facilitate forecasting from past values. In contrast, humidity is strongly influenced by rapidly changing local processes, including convection, pressure perturbations, and precipitation, which introduce high-frequency fluctuations and weak autocorrelations, thereby limiting predictability. Incorporating additional meteorological variables, e.g., pressure and precipitation fields, would likely enhance humidity prediction and enable more comprehensive weather forecasting.
We evaluated support vector regression (SVR) with a radial basis function (RBF) kernel as a nonlinear post–processor for the reservoir readouts. Our implementation employed the SVR module from the scikit-learn library [63], with the regularization parameter $C$ and the tolerance margin $\epsilon$ systematically tuned. FIG. 10 reports the best results obtained. We observe that SVR(RBF) markedly improves temperature prediction for QRC but provides weaker or negligible benefits for ESNs and for humidity prediction. For the difference between QRC and ESNs, we attribute the effect mainly to the dimensionality of the reservoir readout. For QRC (653 features) and ESN(500), the feature dimension remains in a regime where the RBF kernel can provide a meaningful nonlinear mapping and yield observable improvements, with the larger gain of QRC likely related to its quantum-enhanced expressibility. As the ESN size increases to 1000, 5000, or 10000 nodes, the higher dimensionality causes the RBF kernel to suffer from the curse of dimensionality, where distances between different data points become nearly indistinguishable [18]. This limits SVR to extract useful nonlinear structure and increases the risk of overfitting [69, 7], diminishing the benefit of enlarging the network and leading to negligible or even negative gains over linear regression. In the case of humidity prediction, many relevant factors (e.g., pressure and precipitation) are not present in the dataset, which fundamentally limits its predictability and renders nonlinear post-processing much less effective. These interpretations are consistent with the observed behavior, while a more complete understanding may require further investigation.
@@
Appendix I Effects of $T_{1}$ and $T_{2}$ relaxation on the fading memory in quantum reservoir computing
@@
In this section, we examine the influence of the $T_{1}$ and $T_{2}$ relaxation on the fading memory of the quantum reservoir, as quantified by the short-term memory (STM) capacity, through numerical simulations. STM capacity quantifies the ability of the reservoir to reconstruct past inputs from its current state, which is a fundamental property for processing temporal information [29].
@@
First, we reiterate the necessity of incorporating relaxation processes. As mentioned in the main text, an effective quantum reservoir should possess a high-dimensional state space, nonlinear input–output transformations, and fading memory. In unitary systems (commonly employed in quantum algorithm frameworks) the first two properties emerge naturally: the exponential scaling of the Hilbert space provides a high-dimensional representation, while nonlinearity arises from two mechanisms: (i) the input encoding (e.g., rotations yielding sine and cosine responses) [21] and (ii) the multiplicative interaction between sequential inputs [19, 64]. In contrast, fading memory requires that, after a sufficiently long input sequence, the reservoir state converges to a unique trajectory independent of its initial conditions. Due to the distance-preserving nature of unitary dynamics, nonunitary processes (such as $T_{1}$ and $T_{2}$ relaxation) are essential for gradually erasing old information.
@@
Table 3: Parameters of the three-spin diethyl fluoromalonate sample. The left portion presents chemical shifts (diagonal elements, given in the rotating frame) and $J$-couplings (off-diagonal elements), while the right lists the corresponding relaxation times.
In our numerical simulations, we employ a three‑spin system with parameters drawn from diethyl fluoromalonate, a widely used NMR molecular sample, as detailed in Table 3. The spin polarizations are set to $1,\,3.98,\,3.74]\times 10^{-4}$ for carbon, proton, and fluorine respectively. Input is injected into all three spins via the $x$-axis pulse. Considering that the input interval affects the memory capacity, evolution duration $\tau$ is varied from $10^{-4}~{\rm s}$ to $10^{1}~{\rm s}$ to explore its influence. To avoid the uncertainty introduced by selecting partial measurements, the readout is obtained by measuring $4^{3}-1=63$ Pauli operators to capture the full state information. These expectation values are then recalibrated by dividing by the polarization and further perturbed by adding Gaussian noise with $\sigma=10^{-5}$ to improve the robustness of linear regression. For the dataset, we generate a prolonged random sequence $\{s_{k}\}$ uniformly distributed in $[0,1]$. The training and test sets are taken from the tail of the sequence, with sizes of 3000 and 1000 points, respectively. A washout period of size $\max\left\{\left\lceil 10/\tau\right\rceil,t_{d}^{\max}\right\}$ is applied before the training data to eliminate the effects of the initial state, lasting approximately 10 seconds.
@@
FIG. 12(a) illustrates the total STM capacity $C_{\rm STM}$ for various configurations of input intervals and relaxation mechanisms. As expected, $C_{\rm STM}$ vanishes under unitary evolution due to the absence of fading memory. In a unitary system, the reservoir state retains the full influence of the initial state and all prior inputs, preventing the selective extraction of relevant recent information. However, although the introduction of $T_{2}$ relaxation (phase damping) leads to some information erasure, it alone is insufficient to achieve meaningful fading memory. Under unitary dynamics, the quantum system continuously interconverts population and coherence, while dephasing gradually eliminates coherence. This interplay ultimately depolarizes the reservoir, driving it toward the trivial maximally mixed state $\mathds{1}/2^{N}$.
@@
Thus, not every dissipative process is suitable for quantum reservoir computing; the key requirement is a process that drives the system toward a nontrivial state—a role fulfilled by $T_{1}$ relaxation (generalized amplitude damping). Several studies have underscored the necessity of non-unital channels (i.e., channels that do not preserve the identity) in quantum reservoir computing [53, 55]. From an information-theoretic perspective, we conjecture that entropy-decreasing processes are essential for achieving effective fading memory. In amplitude damping channels, the system is driven toward an equilibrium state with lower entropy than the maximally mixed state. Moreover, previous schemes employing input qubit reset [19] and probabilistic state reset [12] incorporate entropy-reducing state preparation processes. In contrast, unitary operations preserve entropy, whereas dephasing increases it [61].
@@
The configuration incorporating $T_{1}$ relaxation exhibits a nonzero STM capacity, with $C_{\rm STM}$ initially increasing and then decreasing as the input interval varies. FIG. 12(b) and (c) show the corresponding behavior of $C(t_{d})$ as a function of delay. In the long-time limit ($\tau\gg T_{1}$), the reservoir relaxes to an equilibrium state independent of the inputs, causing $C_{\rm STM}$ to vanish. For relatively large $\tau$, relaxation erases most of the input history while only retaining the most recent inputs, manifested as a narrow plateau ($C(t_{d})\approx 1$) followed by a rapid decline in $C(t_{d})$. As $\tau$ decreases, the reservoir incorporates information from more input steps; however, nonlinear mixing of these inputs can reduce the effectiveness of linear memory retrieval, resulting in an initial $C(0)<1$ and a gradual decay to zero at larger $t_{d}$. With further decrease in $\tau$, the overall values of $C(t_{d})$ diminish, and in the short-$\tau$ limit, the system behaves similarly to a unitary system, lacking effective fading memory. These results suggest the existence of an optimal input interval that balances the number of retained inputs with high memory fidelity at each delay.